How do you write an equation for the following problems:
1. Kevin weighs 10 pounds more than Tim. The sum of their weight is 470 pounds. How much does Tim weigh?
2. Bryson's new computer costs $178.50 less than his old one. The sum of the costs of the two computers is $3,500. How much did his old computer cost?
3.Kelsey is 4 years older than Kierra. Five years ago the sum of their ages was 14. How old is Kelsey?
4. A rectangle has length three times its width. The perimeter of the rectangle is 56 inches. Find the length of the rectangle.

Answer by ptaylor(1894)

1.
Let x=Tim's weight
Then (x+10)=Kevin's weight
And x+(x+10)=470 get rid of parens
x+x+10=470 collect like terms
2x+10=470
2.
Let x= cost of his old one
Then (x-$178.50)= cost of his new one
And x+(x-$178.50)=$3500 get rid of parens
x+x-$178.50=$3500
2x-$178.50=$3500
3.
Let x=Kelsey's age
(x-5)=Kelsey's age five years ago
Then (x-4)=Kierra's age
((x-4)-5)=Kierra's age five years ago
And (x-5)+((x-4)-5)=14 get rid of parens
x-5+x-4-5=14 collect like terms
2x-14=14

4.
Let x=length
Then x/3=width
Perimeter=2l+2w
And 2x+2(x/3)=56 get rid of parens
2x+2x/3=56 Multiply each term by 3
6x+2x=168

Hope this helps---ptaylor